OpenStudy (anonymous):

Please help me write the function in vertex form and identify its vertex. g(x) = x2 - 10x + 11

OpenStudy (anonymous):

−b2a=102=5 and g(5)=52−10×5+11=25−50+11=−25+11=−14 so vertex is (5,−14) and vertex form will be g(x)=(x−5)2−14

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrvertx.htm this is a good tutorial on vertex form. Sorry, I can't help much more tonight, but I have a calculus test tomorrow that I have to study for. I am sure somebody will check your work. Just use Khan academy or purplemath for tutorials. Khan academy has detailed videos and is free.

OpenStudy (jdoe0001):

hmmmm

OpenStudy (anonymous):

Okay, thank you sm for your hep today! Good luck on your test tomorrow

OpenStudy (jdoe0001):

ahemm @SahiraAlvarado do you know how to "complete the square"?

OpenStudy (anonymous):

No problem. I'm sure I will see you around open study. Nice meeting you.

OpenStudy (anonymous):

You certainly will! I'll be starting my Algebra 2B online class shortly... And it was a pressure meeting you and thanks again for your assistance! It was extremely beneficial.

OpenStudy (jdoe0001):

\(\bf x^2 - 10x + 11\implies (x^2 - 10x) + 11\implies (x^2 - 10x+{\color{red}{ \square }}^2) + 11\) what do you think we need there, in the group, to get a "perfect square trinomial" ?

OpenStudy (anonymous):

25? @jdoe0001

OpenStudy (jdoe0001):

well, yes, \(\bf 5^2 \implies 25\) , keep in mind that all we're doing is borrowing from Zero 0, so if we ADD 25, we have to also SUBTRACT 25, as well so \(\bf x^2 - 10x + 11\implies (x^2 - 10x) + 11\implies (x^2 - 10x+{\color{red}{ 5 }}^2) + 11-{\color{red}{ 5 }}^2 \\ \quad \\ (x-5)^2+11-25\implies g(x)=(x-5)^2-14\)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} g(x)=(x-&5)^2&-14\\ &\uparrow &\uparrow \\ \end{array}\bf \qquad vertex\implies (5,-14)\)

OpenStudy (anonymous):

I got it!!! Thanks